OpenStudy (anonymous):

Assume a stop sign is a regular octagon. A side measures 12.5 in. and a radius is 16.25 in.What is the actual area of the surface of the stop sign?

OpenStudy (radar):

Area= 1/2 (8*12.5)(16.25) Area =(50)(16.25)=812.5 sq in.

OpenStudy (radar):

Do you have an option with that value?

OpenStudy (radar):

Area = (1/2) X perimeter X Apotherm Where apotherm is the distance from the center to the bisector of a side.